GreenHab Report – December 11th

  

Crew 217 GreenHab Report - 11-12-2019

GreenHab Officer: Cynthia Hori

Environmental control:

Heating

Shade cloth (40%) on

Average temperatures

Low 66 F

High 90 F

Hours of supplemental light :5

Daily water usage for crops: 80 L

Daily water usage for research and/or other purposes:

Water in Blue Tank— 300 gallons

Time(s) of watering for crops: 7:30 pm

Changes to crops: Pepper seedlings emerging growing quite well. Paper
whites were rotated from top to bottom shelf. Getting more water and
growing quite well

Narrative: We didn’t get our work done today that we said we would.
We have the Greenhab slated for tomorrow. A second shade cloth (30%)
will be added tomorrow and squash will be transplanted. New lettuce
seeds going in tomorrow. The herb garden is getting tall and the
oregano is beginning to lift off of the soil and gain about 2 cm in
height. Radishes may need to be thinned more.

Harvest: none

Command Report – December 11th

Commander’s Report

11 Dec 2019

Shannon Rupert

It seems strangely familiar. Although I haven’t suited up for an EVA
in over a decade, the routine is the same. Slide flight suit on.
Well, this one doesn’t quite fit the way the last one I had put on
did. Instead of slipping into the oversized suits of the past, I pull
and tug and grunt as I squirm my way into a flight suit one size too
small. It was either this or steal the larger one from Atila, and
pride keeps me from doing that. Once the flight suit is on, I add the
radio. Earpiece, check. Radio with good battery, check. COMMS
check, check. I am ready. I chose the prototype for the one piece
suits, sometimes called the milk carton suits, for this EVA. It takes
three people to adjust it to fit me, because one shoulder strap,
incorrectly threaded, just keeps popping over and over. I can’t see
them, but I can feel them as they work on the suit. Finally, I feel
the strap tighten and stay that way, and I know someone has found a
way to fix it. Thumbs up. I’m ready now. Into the airlock. This too
seems familiar, something I am comfortable with. We wait until the
light goes off and are given permission to exit. I open the heavy
door and lift one foot, looking down. For one panicked second, I
hesitate, because this now seems strange. Whether from age or from
lack of practice, I worry that I won’t be able to walk. But I do. I
get to the rover and bump my suit a few times on the frame as I settle
in. I find myself in a reclining position as Jen starts the drive to
our EVA destination. I laugh as I sit there, getting flung from side
to side like a damn ragdoll. I spend the next twenty minutes trying
to get comfortable. I ignore the incredible Mars-like landscape we are
passing through. That I know like the back of my hand. What I don’t
know is how to get comfortable in this damn suit. Finally, I lean
forward and realize that there is a way to sit up straight. But it
comes at a price. When I sit up, the helmet pummels the hell out of
me. First my lip, then my teeth, then my lip again. And all the time
I am laughing, because what the hell do you do in that situation?
Finally, after what seems the longest rover ride of my life, we arrive
at the Moon overlook. I realize with delight that I am small enough to
just slither right out of the rover seat, and I’ve got my feet on
solid ground. We take a few fun photos, and then I head up the road to
scout for gypsum. I find a promising mound, then stop. I radio my
team “Do you think if I get down on my knees I will ever be able to
get up again?” I plop onto my knees and remember in a flash how hard
it is to collect samples when your vision is limited by the helmet and
your gloves make it difficult to pick things up. We gather some
gypsum and my team holds my hand so I can get up. We have two more
sampling sites, and as we head to the next one, I wonder why the sun
is getting so low in the sky. At the final sampling site I realize
that I had screwed up a long held rule at MDRS and I have scheduled
our EVA to end at sunset and not one half hour before sunset. That’s
what Atila had been trying to tell me, when I insisted the EVA go
until 5. We travel back to MDRS at a fast pace. Jen is enjoying the
hell out of driving the rover, but I am distracted by the way my
helmet keeps trying to punch me in the face. We park, enter the
airlock and in a few minutes I’ve put the murderous suit back on the
shelf. Both it, and I, will live to fight another day.

Journalist Report – December 11th

  

Crew 217 Journalist Report
11 Dec 2019
Author: Teresa Hislop

“Failure is not an option” was today’s theme.

There were no petri dishes in the science dome. Our research plans
call for plating samples from two separate collections: gypsum
samples collected at Moon Overlook and green sediment samples from the
Morrison strata in Lith Canyon. Not be be deterred, we made “petri
bags” by mixing agar and pouring it into zip log bags. Success!

EVA One (Ann, Cynthia, Kevin, Teresa) journeyed to Lith Canyon. One
of MDRS’s absolute rules is that one stops when the Rovers’ battery
reaches 60%. Do not pass go. Do not collect $200. Do not take the
Rover further afield when the battery reads 60%. About two-thirds
the way to Lith Canyon, Spirit’s battery read 60%. Not to be deterred,
the astronauts continued with their mission, one Rover down. Failure
was not an option.

They successfully collected water samples from three separate
potholes. They also took measured pH and temperature and took
salinity readings They also scaped rock samples from four samples from
the green sediment in the Morrison strata: one from above the green
layer, one from the green layer, one from under the green layer, and
one from a group of crystalline structures in the layer above the
green layer. Finding themselves with over 60 minutes left in their
EVA allotment of time, the group also collected gypsum samples at the
Cow Dung Road/Moon Overlook Road junction.

While waiting for the radios and Rovers to recharge the group ground
and plated the Morrison strata and gypsum samples. The gypsum
samples were plated on 20% saline agar; the Morrison strata samples
were plated on nutrient agar and incubated.

EVA Two drove to Moon Overlook, stopping at three different locations
to collect gypsum. They went out and back without incident but not
without adventure. It was the first time in over 10 years Shannon had
been in a space suit and the first time ever that she has ridden in a
Rover in a space suit.

Lesson and idea sharing continues. Today’s highlight is Jen’s
Flipgrid app. She posts videos that her students view and respond to.
She can see and respond to their responses. It is a fabulous way for
asynchronous communication from “Mars” to Earth and back again. 

Science Report – December 11th

  

Research Report 11 December 2019

Crew 217 – NSBU 217

Sol 2

Submitted by Science Officers Kevin Berean & Ann Rockett

1. Sharing of Mars and Space-related Classroom Activities:

Crew members spent 4 hours on SOL 1 in a routable discussion sharing
Mars and space-related classroom activities. Additional time has also
been spent on SOL 1 and SOL 2 as the SIM unfolds discussing ways in
which we can incorporate this firsthand experience into classroom
lessons and activities.

2. Conducting a Biodiversity Survey of Ephemeral Potholes in a
Mars Analog:

Four crew members conducted an EVA on SOL 2 (that commenced at 10:15AM
and ended at 12:30PM) and biodiversity samples were extracted from
three ephemeral potholes at 4256500, 518500. PH and salinity readings
were also taken and recorded for each pothole. Samples from Pothole 1
and Pothole 2 underwent a preliminary examination under a microscope
in the MDRS Science Dome. Further examination of said samples will
take place on SOL 3.

3. Investigating if Life Exists in a Green Layer of the Morrison
Strata that is Exposed Throughout the MDRS Exploration Area:

Four Crew Members conducted an EVA on SOL 2 (that commenced at 10:15AM
and ended at 12:13PM) and samples were extracted from a green layer of
the Morrison Strata at 4256500, 518500. Three samples were taken for
investigation: sample 1 from the bottom of the green layer; sample 2
form the middle of the green layer; and sample 3 from the top of the
green layer. All three samples were ground, plated and placed in the
incubator at room at 27°C.

4. Search for Halophiles in Surface Gypsum:

Gypsum samples were collected during the two EVAs today. Gypsum
samples from the morning EVA (that commenced at 10:15AM and ended at
12:30PM) were collected on Cow Dung Road and ground and plated. Gypsum
samples from the afternoon EVA (that commenced at 3:40PM and ended at
5:00PM) were collected at the Moon overlook.

5. Exploration of Ways to Communicate with Classrooms at MDRS.

The crew continues to brainstorm ways in which future NSBU crews can
communicate with classrooms during the SIM. Setting up a Flipgrid
classroom is one way NSBU crews can communicate with classrooms during
a SIM. In a Flipgrid platform, the NSBU teacher can post videos from
MDRS during SIM to “the grid” that his/her students can access and
watch via their established Flipgrid accounts. The students can then,
in turn, post questions via video back to the NSBU teacher for his/her
response. Students can also pose questions to each other via the
Flipgrid. This method is currently being piloted and tested by one of
our crew members and she has been able to post a video for her
classrooms on SOL 1 and SOL 2, which has generated 18 question/reply
videos from her students. The only issue that has surfaced using this
method is the difficulty establishing an internet connection. 

Operations Report – December 11th

  

Crew 217Operations Report 12/11/ 2019

SOL: 1

Name of person filing report: Hope Zubrin

Non-nominal systems: radios

Notes on non-nominal systems: radios still having charging issues and
they all have new batteries

Generator:

Hours run: 12

From what time last night: 7:00 pm

To what time this morning: 7:00AM

List any additional daytime hours when the generator was run: n/a

Solar— SOC 48 % (Before generator is run at night)

Diesel Reading – 22%

Station Propane Reading –not checked

Ethanol Free Gasoline: 5 gallons

Water (loft tank): 55 gallons

Water Meter: 1467945 units

Water (static tank): 425 gallons

Static to Loft Pump used – yes

Water in GreenHab: 300 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: off campus

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used:yes

Hours: 118.0

Beginning charge: (Before EVA) 100 %; 100%

Ending charge: (On return from EVA, before recharging) 42 %;38%

Currently charging: Yes

Opportunity rover used: off campus

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: yes

Hours: 123.3

Beginning charge:100%; 100%

Ending charge: 50 %; 52 %

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:not used

Oil Added? n/a

ATV Fuel Used: zero gallons

# Hours the ATVs were used today:

Notes on ATVs: nothing to report

HabCar used and why, where? No

CrewCar used and why, where? In Grand Junction

General notes and comments:

Summary of internet: nominal

Summary of suits and radios: all suits operational see radio notes

Summary of Hab operations: nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Dissection scope needs repair. Many
supplies needed: Petri dishes, sample cups, lab tape, sharpies,
glassware, nutrient agar, scissors, and other consumables.

Summary of RAM operations: nothing to report

Summary of any observatory issues: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: nothing at this time. 

EVA Report – December 11th

  

Crew 217 EVA Request 12-11-19

EVA #: 3

Today’s date: Sol 2

Requested EVA Date: Sol 3, 12-12-19

Requested start time for EVA: 10:00 am

Requested end time for EVA:13:00 pm

Weather is clear and sunny supporting the requested EVA.

Purpose of EVA:

To collect ancient invertebrates for presentation to students upon
return to Earth.

Destination:

Walk from Hab up Sagan Street to Hab Ridge (destination)

Coordinates: 518,000, 4251,300 UTM

EVA Participants: Cynthia, Teresa, Kevin, Ann, Jen, Hope 

Sol Summary – December 11th

  

Crew 217 Sol Summary Report 11Dec2019

Sol: 2

Summary Title: Mars needs teachers

Author’s name: Atila Meszaros

Mission Status: Second day on the red planet and rocking our mission goals.

Sol Activity Summary: We started the day earlier, as we needed to cover a considerable amount of activities. The morning began with a double breakfast on my side and our first Biology class from our favorite Ecology teacher, Shannon Ruppert. The main objective of the day was to get 3 types of biology samples: gypsum from the Moon areas, water from pre-selected ponds and some fragments from the mysterious unidentified green layer. All of them were collected wonderfully. We aimed to plated them on agar so we would be able to determine qualitatively the biological content of them. In order to do that, we MacGyvered the agar making business and created (patent pending) the "Petry baggies". Today has been a wonderful day, I saw Shannon put a suit and go on an EVA for the first time ever, and you don't usually see that. I haven't been in sim for a while, but spending this week with wonderful people, one more inspiring and dedicated than the next one, it's really fulfilling. I haven't been in school for a long long time, but if I had one of these teachers for just a day when I was growing up, I would have definitely changed my life. Looking forward to seeing their work and how this experience will create new and better programs for their kids.

Look Ahead Plan: Last day of the mission! We will start preparing the rockets for our afternoon EVA. Part of the crew will be doing the first last EVA to Hab Ridge during the morning. After lunch, we will be working on the GreenHab, planting and making a terrarium with Mars soil simulant. During the afternoon, small rockets wait for us to be launch And during the evening, we will try to wrap up the mission and spend our last night together.

Anomalies in work: Nothing to report.

Weather: Colder than yesterday, partly clear skies, light winds, and an impressive full moon.

Crew Physical Status: Healthy, a little bit tired, but happy.

EVA: Our first two long EVAs tackling our scientific goals: water holes, gypsum, and unidentified green layer sample return.

Reports to be filed: Sol Summary, EVA Reports 3 & 4, EVA Request, GreenHab Report, Science Report, Operations Report, Journalist Report, and Photos.

Support Requested: Science Dome materials and instruments (listed on the Science Report).